# Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models

### v2

@@ 4.2 Experiments on Llama2-7B
Llama2-7B LoRA 89.9M 38.97 0.44
PiSSA 89.9M 43.52 6.92
Bone b=16 21.7M 45.79 3.54
Bone b=32 43.5M 46.70 7.67
Bone b=64 87.0M 49.36 8.88
Bone b=128 174.0M 53.15 10.56
Figure 4: Training loss curves of Bone with different block_size on the MetaMathQA dataset.
4.4 Ablation experiments on different grouping methods.
In this subsection, we explore the impact of different grouping methods in the Bone structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bone-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  5. Since in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bone structure, making it difficult to align parameter counts. Even though Bone-row performs better than Pissa with only around 20M parameters, the gap between Bone-row and Bone-col is too large. Initially, we suspected this gap was due to Bone-row lacking about 14M parameters compared to Bone-col. However, this assumption was disproved when we observed that Bone-row with 32 blocks and 43.5M parameters performed similarly to Bone-row with 72.8M parameters. This made it difficult to directly evaluate which grouping method performs better.
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bone fine-tuning. This allows for a fairer comparison between Bone-row and Bone-col. The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bone can effectively fit data regardless of the grouping method used. The key factor influencing Bone’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bone will need to adjust its grouping or employ other techniques to adapt to these new technologies.
Table 5: Comparing Bone-row, Bone-col on math tasks
@@
Llama2-7B Bone-row 72.8M 45.76 7.82
Bone-col 87.0M 49.36 8.88
RWKV6-3B Bone-row 55.1M 25.93 3.32
Bone-col 55.1M 25.25 2.82
4.5 Ablation experiments on Block Affine.
We will conduct ablation experiments on the Block Affine structure, which consist of two parts:
1. Verifying the effectiveness of “Weight Guide” on LLaMA2-7B: The results are shown in the table  6. It is evident that the performance of Bone-Unconstrained is even worse than that of Bone-Col $b=32$, which has half the number of parameters but incorporates the “Weight Guide” characteristic.
2. Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
The comparison formulas are as follows:
$Bone-Unconstrained\Delta W_{n,n}=W_{n,n}A_{n,n}+B_{n,n}$ (11)
$Bone-Hadamard\Delta W_{n,n}=W_{n,n}*bone_{n,n}+bone_{n,n}$ (12)
Table 6: Block Affine on math tasks
@@
Llama2-7B Bone-col b=32 43.5M 46.70 7.67
Bone-row b=64 72.8M 45.76 7.82
Bone-Unconstrained b=32 72.8M 44.35 6.16
RWKV6-3B Bone 55.1M 25.93 3.32
Bone-Hadamard 55.1M 22.44 2.67
@@ 5 Resource and efficiency
Table  7 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU. The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512. The structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage. To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput. At the end of the table, we provide the real-world training scenarios comparing the memory usage and token throughput of Bone and LoRA.
Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
@@
RWKV-1.6B LoRA 33.0M 7314 MB 5.86 kt/s
Bone 33.0M 12276 MB 4.11 kt/s
Bone(checkpoint) 33.0M 7398 MB 3.46 kt/s
Bone(model checkpoint) 33.0M 7398 MB 1.59 kt/s
@@
Bone 55.1M 22978 MB 2.16 kt/s
Bone(checkpoint) 55.1M 12168 MB 1.72 kt/s
RWKV-3B LoRA(model checkpoint) 55.1M 4*15328 MB 15.6 kt/s
Bone(model checkpoint) 55.1M 4*15305 MB 14.2 kt/s
@@ 6 Conclusion
This paper introduces a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information. Based on this theory, we designed a new PEFT technique called Bone (Block Affine). Additionally, we employed block matrix multiplication to efficiently promote information fusion within the weights in the Bone structure. Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA. Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility. Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
7 Limitations and future work
Due to limited training resources, this paper lacks comparative experiments on full fine-tuning. However, the extensive experiments presented demonstrate that Bone possesses strong data fitting and generalization capabilities, paving the way for new directions in PEFT development. However, there are still some issues that require further consideration and resolution:
1. While the Bone architecture shows better training performance, its peak memory usage is higher compared to LoRA and its variants.
2. It remains to be seen whether Bone is suitable for large models in other modalities.

### v4

@@ 4.4 Bone vs Bat
Llama2-7B Bone 87.0M 48.16 8.58
Bat 87.0M 49.36 8.88
RWKV6-3B Bone 55.1M 41.73 6.52
Bat 55.1M 42.76 6.60
4.5 Ablation experiments on different grouping methods for Bat
In this subsection, we explore the impact of different grouping methods in the Bat structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bat-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  5. Since in LLaMA2-7B, the dimension of gate_proj in the MLP part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bat structure, making it difficult to align parameter counts. Although Bat-row uses 15M fewer parameters than Bat-col, it still delivers excellent performance. However, this discrepancy in parameter counts makes it challenging to accurately evaluate the differences between the grouping methods.
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bat fine-tuning. This allows for a fairer comparison between Bat-row and Bat-col. The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bat can effectively fit data regardless of the grouping method used. The key factor influencing Bat’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
Table 5: Comparing Bat-row, Bat-col on math tasks
@@ 4.5 Ablation experiments on different grouping methods for Bat
Llama2-7B Bat-row 72.8M 45.76 7.82
Bat-col 87.0M 49.36 8.88
RWKV6-3B Bat-row 55.1M 25.93 3.12
Bat-col 55.1M 25.25 3.09
4.6 Ablation experiments on the specific implementation of “Weight Guide”
Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
The formula for Bone-Hadamard is as follows:
$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (5)
$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\odot bone_{d/b,b,b}+bone_{d/b,b,b})$
Table 6: Block-Affine on math tasks
@@ 4.6 Ablation experiments on the specific implementation of “Weight Guide”
RWKV6-3B Bat 55.1M 25.93 3.32
Bat-Hadamard 55.1M 22.44 2.67
@@ 5 Resource and efficiency
Table  7 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU. The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
The results show that Bone has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient. However, Bat incurs significantly higher memory usage due to large intermediate values and is slower in comparison.
At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
Therefore, a key focus of our future work will be improving the Bat operator to enhance token throughput and reduce memory usage.
@@
Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.95}$ kt/s
Bat 55.1M 22978 MB 2.16 kt/s
RWKV-3B LoRA(use checkpoint) 55.1M 4*15328 MB 15.6 kt/s
Bone(use checkpoint) 55.1M $\mathbf{4*15304}$ MB $\mathbf{16.0}$ kt/s
Bat(use checkpoint) 55.1M 4*15305 MB 14.2 kt/s
@@ 6 Conclusion
This work, inspired by GQA and MQA, leverages the sparsity of LLM weights to design the Block-Affine Adaptation (Bone) structure. In Bone, the original weights are divided into multiple subspaces, all of which share a single low-rank matrix initialized to zero for updates. Extensive experiments demonstrate that Bone consistently outperforms LoRA and its variants across various tasks, while also offering superior computational efficiency. To break the limitations of LoRA and effectively utilize the information from the original weights, we propose the “Weight Guide” theory, which enables significant improvements through simple linear transformations.
By integrating these innovations, we introduce a new structure called Block-Affine Transformation (Bat). Bat not only validates the effectiveness of “Weight Guide” but also addresses the limitation of Bone, where identical updates are applied to all subspaces. Experimental results show that Bat surpasses Bone on multiple tasks, although with a trade-off in computational efficiency.
Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
7 Future work
Bone and “Weight Guide” are merely a starting point, offering a foundation for researchers to explore additional branches inspired by Bat. There are many directions for future work.
1. Can Bone adapt to multimodal tasks and other complex scenarios?
2. Can “Weight Guide” be implemented using alternative computation methods?